A 5kg block , resting on a rough horizontal surface is connected by a light inextension string passing over a light frictionless pulley to a second block of mass over a 3kg hanging vertically . Applied force F acting on the 5kg block and the coefficient of kinetic friction is 0,2 . Calculate the magnitude of the vertical component of F if the magnitude of the horizontal component of F equals 38 N


1
Expert's answer
2021-04-06T13:51:39-0400
"N=Mg-F_y\\F_f=\mu N=\mu(Mg-F_y)\\ma=mg-F_f-F_x=0\\mg=\mu(Mg-F_y)+F_x
\\(3)(9.8)=0.2((5)(9.8)-F_y)+38\\F_y=92\ N"